OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? 2x+y=-8 x-y=-4

OpenStudy (anonymous):

@Spring98

OpenStudy (spring98):

So first solve for x and y add the 2 equations together 2x+x+y-y=-8-4 3x=-12 divide by 3 x=-4 subsitute -4-y=-4 addd 4 to both sdies -y=0 the solution is (x,y) (-4,0) or D